OpenStudy (kathy_morales13):

Help please ..............Catie is 3 years older than her brother. The sum of their ages is no more than 31.

OpenStudy (ben00):

How old is the brother?

OpenStudy (danjs):

Call 'C' Catie age, and 'B' the brothers age... Catie is 3 years older than her Brother... C = B + 3 The sum of their ages is no more than 31... \[C + B \le 31\]

OpenStudy (danjs):

Put the first one in for C in the second. \[B + 3 + B \le 31\] \[B \le 14\] The Brother B is less than or equal to 14, so Catie is less than or equal to 14 + 3 = 17

OpenStudy (kathy_morales13):

okay thank you :)